OpenStudy (anonymous):

Find the complex zeros of the polynomial function. Write f in factored form. f(x)=x^3-9x^2+28x-30

OpenStudy (anonymous):

Do you know how to start this?

OpenStudy (anonymous):

I am stuck actually.

OpenStudy (anonymous):

How much work have you done?

OpenStudy (anonymous):

Did you use the quadratic formula?

OpenStudy (anonymous):

If I were to solve this, I would first factor, and then solve the remaining part by using the quadratic formula.

OpenStudy (anonymous):

So, the factored form looks like:

OpenStudy (anonymous):

(x-3)(x^2-6x+10)=0

OpenStudy (anonymous):

Do you follow so far?

OpenStudy (anonymous):

yes I got that part

OpenStudy (anonymous):

okay, now solve each part (x-3)=0 (x^2-6x+10)=0.....solve this using the quadratic formula. You'll get two imaginary roots.

OpenStudy (anonymous):

Or you can solve: (x^2-6x+10)=0..... by completing the square.

OpenStudy (anonymous):

okay so that part is x=6+2i/2, x=6-2i/2

OpenStudy (anonymous):

Yes, you can also further simplify those imaginary roots and get: x=3+i and x=3-i

OpenStudy (anonymous):

So you have x=3, x=3-i, x=3+1

OpenStudy (anonymous):

That pretty much means that there is a triple root at x=3

OpenStudy (anonymous):

so x=3 is what it should be written as or x=3, x=3-i, x=3+1

OpenStudy (anonymous):

It should be written as x=3, 3+/-i...... +/- means plus over minus.

OpenStudy (anonymous):

ok I see, thanks.